Hi, how do you solve the following word problem?
Arielle has a collection of grasshoppers and crickets, She has 561 insects in all. The number of grasshoppers is twice the number of crickets. Find the number of each type of insect that she has.

Answer by mananth(16946)

let the number of crickets be x
grasshoppers will be 2x
Total insects =561
..
x+2x=561
3x=561
x=561/3
x=187
crickets = 187
grasshoppers = 2x = 2*187= 374